kekeman:

At which values in the interval [0, 2π) will the functions f (x) = cos 2x - 5 and g(x) = sin3x - 5 intersect? A. x = 0, π/10, π/2, 9π/10, 13π/10 B. x = 3π/10, π/2, 7π/10, 13π/10, 17π/10 C. x = π/10, π/2, 9π/10, 13π/10, 17π/10 D. x = 0, π/2, π/10, 13π/10, 17π/10

Narad:

Intersection will be when \[f(x)=g(x)\] That is \[\cos2x-5=\sin3x-5\] \[\cos2x=\sin3x\] \[\cos2x=\cos(\frac{ \pi }{ 2 }-3x)\] \[2x=\frac{ \pi }{ 2 }-3x +2k \pi\] and \[2x=-\frac{ \pi }{ 2 }+3x+2k \pi \] Where \[k \in \mathbb{Z}\] By giving different values to k, you will get the right solution

kekeman:

So what is the answer?

kekeman:

@surjithayer

imqwerty:

Set the limits on x to be between 0 and \(2\pi\). That'll give you a range for k, k can take all the integer values in this range. Plug in those values of k, and that'll give you your x

kekeman:

I don't know how to